OpenStudy (anonymous):

Give 3 examples of simplifying Radicals. Exaplain and show work. Solve these examples also: Simplify The square root of 28. The square root of 132. :) thanks for the help;)

OpenStudy (anonymous):

Any help?

jhonyy9 (jhonyy9):

sqrt28 =sqrt(4*7)=sqrt(2^2 *7) =2sqrt7 do you understand it sure ?

jhonyy9 (jhonyy9):

sqrt132 = ?

OpenStudy (anonymous):

I don't understand it thr much. Can you explain alittle more please

jhonyy9 (jhonyy9):

how can you rerite the 132 using perfect square number ?

OpenStudy (anonymous):

I don't know

jhonyy9 (jhonyy9):

so the same like 4 here in case of number 132 than 132/4 =33 than you will can rewrite sqrt132 = sqrt(4*33) = sqrt(2^2 *33) = 2sqrt33 this is because you know that sqrt of a perfect square number is this number like sqrt9 = sqrt(3*3) = 3 or sqrt16 = sqrt(4*4) =4 or sqrt121 = sqrt(11*11) = 11

jhonyy9 (jhonyy9):

or for example sqrt(x^3) = sqrt(x^2 *x) = xsqrtx ok ?

OpenStudy (anonymous):

Ok

jhonyy9 (jhonyy9):

so than good luck bye

OpenStudy (anonymous):

Wait so 132 squared is 2 x33 squared ? Simplified?

jhonyy9 (jhonyy9):

no no squareroot of 132 will be simplifie it in the form 2 time squareroot 33 because 132 can be writing like 4*33 and you know that 4 is a perfect square of 2

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

Have a nice day ;)

jhonyy9 (jhonyy9):

so than result that you have understood it right all good luck bye